OpenStudy (anonymous):

Find a simple basic function as a right/left end end behavior model f(x)=e^x-2x

OpenStudy (dan815):

okay

OpenStudy (dan815):

u get this one already??

OpenStudy (dan815):

y = e^x - 2x

OpenStudy (anonymous):

i need help with this one i dk wht to do

OpenStudy (anonymous):

do you know what to do ?

OpenStudy (dan815):

think about e^x

OpenStudy (dan815):

for larger values of e^x does the linear line 2x values really matter

OpenStudy (anonymous):

no

OpenStudy (dan815):

but for small values of x e^x - x will start to matter and for e^x where x is negative

OpenStudy (dan815):

e^x will get smaller and smaller while 2(-x) will stay about the same

OpenStudy (anonymous):

ok? but were does this all lead to the point about left n right end?

OpenStudy (dan815):

here u can probably benefit from this, just dont rely on this for all your answers try to understand it!

OpenStudy (dan815):

http://www.wolframalpha.com/input/?i=e%5Ex-2x

OpenStudy (anonymous):

yea that is what i want to do learn it

OpenStudy (dan815):

im basically saying for values of x as you go right the e^x is dominant and as you go toward the left, the -2x is dominant

OpenStudy (anonymous):

oo i get that part now but what does simple basic form

OpenStudy (anonymous):

how do you get that

OpenStudy (dan815):

as really big values of x, ull only see the effects of e^x at really small values(more and more negative) of x, you will only see only -2x as e^-big number = approaching 0

OpenStudy (dan815):

e^-inf = 1/e^inf = 0

OpenStudy (dan815):